# Hijacking Context in Large Multi-modal Models

### v1

@@ Abstract
Recently, Large Multi-modal Models (LMMs) have demonstrated their ability to understand the visual contents of images given the instructions regarding the images. Built upon the Large Language Models (LLMs), LMMs also inherit their abilities and characteristics such as in-context learning where a coherent sequence of images and texts are given as the input prompt. However, we identify a new limitation of off-the-shelf LMMs where a small fraction of incoherent images or text descriptions mislead LMMs to only generate biased output about the hijacked context, not the originally intended context. To address this, we propose a pre-filtering method that removes irrelevant contexts via GPT-4V, based on its robustness towards distribution shift within the contexts. We further investigate whether replacing the hijacked visual and textual contexts with the correlated ones via GPT-4V and text-to-image models can help yield coherent responses.
@@ 1 Introduction
Large Language Models (LLMs) pre-trained with huge amount of text corpus have demonstrated remarkable generalization on various language tasks such as human-like dialogues google_2023 ; openai2023 ; openai2023gpt4 , reasoning-tasks wei2021finetuned ; lewkowycz2022solving ; yao2022react , and few-shot in-context learning min2021metaicl ; min2022rethinking where a few demonstration examples are given as a context for answering the test question. By bootstrapping LLMs with several add-on layers, Large Multi-modal Models (LMMs) openai2023gpt4 ; fromage ; llava ; minigpt were enabled to understand the visual contents and answer the corresponding textual instruction. LMMs also inherited the in-context learning capability of LLM, where a coherent sequence of images and textual information are given as a context for the input prompt fromage ; openai2023gpt4 .
In this paper, we identify a new limitation of these off-the-shelf LMMs, hijacking contexts, where a small fraction of incoherent images or text descriptions mislead them to only generate biased output about the incoherent contexts. In Figure 1, we provide the qualitative example of context hijacking on visual story telling (VIST) dataset vist with FROMAGe fromage as LMM. Given all the former sequence of images and their corresponding captions, the LMM is asked to provide a caption or description of the final query image. Normally, when all the visual and textual information is coherent and consistently aligned, LMM successfully provides a coherent response under consideration in the given contexts. However, when only a single image-text pair that contains an irrelevant subject is appended to the context, LMM falls into the hijacked context and loses coherency concerning the bunch of formerly given original contexts. In real-world applications and scenarios of LMMs, there is no guarantee for the inexistence of such noises and irrelevant contexts. For reliable usage of LMM, it should be robust to such distribution shift and stick to the majority of context without confusion on the minority of irrelevant contexts.
@@
To address this, we propose a pre-filtering method that removes irrelevant and hijacked contexts via prompting into GPT-4V openai2023gpt4vision , based on its robustness towards distribution shift within the contexts. Notably, we observed that GPT-4 and GPT-4V have the capability to filter irrelevant textual or visual information via a simple few-shot prompting, respectively. Our goal is to leverage this remarkable capability to enhance off-the-shelf LMMs that have a smaller amount of parameters and are pre-trained on a much smaller scale of dataset compared to GPT-4 and GPT-4V.
One step further, we investigate whether modifying the hijacked image and text contexts with the correlated ones can help yield better coherent responses. Firstly, we instructed GPT-4 to identify any irrelevant sentences and replace them with appropriate alternative sentences that convey coherency given in the contexts. Subsequently, we prompted text-to-image diffusion models  [^1] We employed DALLE-3 dalle3 for its detailed sensitiveness to the input prompts to generate a coherent image corresponding to the newly replaced sentence under consideration for the original sequence of sentences and images. As a result, a new alternative image and caption that are coherent with the original story plot are interleaved into inputs for LMMs. Notably, we observed that LMMs still exhibit hijacked responses, due to two possible reasons: 1) the generated captions are still confusing which seem to be the combination and mixture between the original and hijacked context. 2) the generated images are not realistic and do not have the consistent texture or styles with the original image sequences.
We showcase our problem statement and aforementioned methods and investigations with qualitative examples one by one [^2] We tested with FROMAGe fromage as the baseline of LMM.. In summary, our key contributions are outlined as follows:
@@ 2 Related Works
Recently, large language models (LLMs) have gained considerable versatility based on enlarging the scale of pre-training datasets gpt-2 ; gpt-3 ; openai2023gpt4 . Along with the development of human-aligned feedbacks rlhf , LLMs have further been well-suited for conforming to the instruction rather than predicting simple meaningless next tokens and words. These LLMs have also demonstrated versatility on various language domain tasks including in-context learning gpt-3 ; chan2022data and long-term dialogues jozefowicz2016exploring ; dai2019transformer ; liu2023lost .
@@
Despite the compelling performance of LLMs on language tasks, LLMs could not understand the visual contexts in an image. To bootstrap LLMs with vision modality, several works llava ; minigpt ; fromage employed an additional pre-trained visual encoder and leveraged a few linear layers to adapt the visual tokens to be understandable by the frozen LLMs. FROMAGe fromage further demonstrated the in-context learning capability of LMMs where consecutive sequences of image-text pairs are given as a context to respond to the query image. In this paper, we focus on the vulnerability of LMMs to incoherent and irrelevant context information.
@@
In the language domain, several works empirically observed that LLMs are easily hijacked by irrelevant context information shi2023large , small perturbations of characters ($e.g.,$ from "film" to "fi1m") wang2023adversarial , and adversarial addition of suffix words qiang2023hijacking . In the vision domain, there have been several approaches that investigated the adversarial robustness of LMMs schlarmann2023adversarial ; qi2023visual where a small adversarial tweak of pixel values significantly misled LMMs to output violent content. To the best of our knowledge, our work is the first to investigate hijacking contexts in an in-context learning scenario where the consecutive sequence of images and texts along with irrelevant ones are given as the context to provide answers about the query image.
3 Preliminaries
Large Language Models.
An autoregressive LLM $p_{\theta}$ is pre-trained on a large corpus of text tokens with the next-token prediction objective. Formally, given a text of image caption $y$ and its sequence of tokens ($\bm{s}=s_{1},...,s_{t-1}$) separated by a BPE tokenizer sennrich2015neural , the model is trained to maximize log-likelihood for the next token $s_{t}$:
@@ 3 Preliminaries
Large Multi-modal Models.
@@
For in-context learning scenarios where multiple images and corresponding text descriptions are given, FROMAGe fromage is trained with a random concatenation strategy where multiple image-text pairs are piled for attending and responding more explicitly to each image. Hence, FROMAGe exhibited superior capability for answering the query image $x_{n}$ given former visual contexts ($x_{1},\ldots,x_{n-1}$) and textual contexts $(\bm{s}_{1},\ldots,\bm{s}_{n-1})$:
@@
4 Method
Filtering hijacked contexts.
To address the context hijacking issue, we first propose a simple yet effective method that can preemptively remove hijacked visual or textual contexts. We observed that GPT-4 and GPT-4V have the capability to filter irrelevant textual or visual contexts via a simple few-shot prompting, based on the powerful zero-shot performance and its robustness towards distribution shift within the contexts. As in Figure 2, we labeled the sequence number for each image for GPT-4 to understand both the image and text pair correspondingly. Then, we prompted GPT-4 to remove any irrelevant images or textual contexts based on the story plot. GPT-4 appropriately identifies the hijacked contexts (sequence number 5) and hence these filtered outputs will prevent LMMs (such as FROMAGe) from yielding hijacked responses.
Figure 2: Filtering hijacked contexts via GPT-4(V). Given multiple image-text pairs of a visual story plot from VIST dataset vist , GPT-4 is instructed to filter any irrelevant images or text descriptions based on the context.
Reforming hijacked contexts
We further investigate whether reforming the identified irrelevant visual and textual contexts into coherent ones can help LMMs yield better coherent responses. As in Figure 3, to reform the hijacked text description, we first instructed GPT-4 to provide appropriate sentences that convey coherency given in the contexts. Subsequently, we instructed DALLE-3 to generate a synthetic image that corresponds to the newly reformed text, with the consideration of underlying visual and textual context information. Consequently, the arranged sequence of visual and textual contexts with the newly reformed one might help to provide a more sufficient and coherent context for LMMs not to be confused.
Figure 3: Reforming hijacked text and images via GPT-4 and DALLE-3. Given multiple image-text pairs of a visual story plot from VIST dataset vist , GPT-4 is instructed to replace any irrelevant images or text descriptions with coherent ones. Subsequently, DALLE-3 is instructed to generate an image corresponding to the newly reformed text, under consideration of the underlying context.
@@ 5 Discussions
We investigate the effect of the sequential location of hijacked context on the coherency of the response from LMMs. In Figure 4, when the hijacking context is interleaved in the former part, the final response still tends to adhere to the original context, "family". In contrast, as the hijacking context is interleaved closer to the query image, the response tends to be incoherent to the former contexts and biased towards the hijacked context, "football".
@@
In Figure 5, we observed that our reforming strategy introduced in Section 4 still does not ensure LMMs to give coherent responses. We conjecture that the reformed captions still contain the hijacked context ("football") while also adhering to the original context ("family"). The generated image also correspondingly contains both visual contexts of "family" and "football", confusing LMMs to produce biased responses. Also, we conjecture that the style and texture of the reformed image differ to the original images, which might also induce LMMs to be confused.
@@ 7 Conclusion
In conclusion, we newly identified the context-hijacking problem of existing LMM where the irrelevant visual and textual contexts induce the LMM to generate a biased and incoherent response, disregarding the original context. To mitigate this issue, we introduced a simple filtering technique to preemptively remove such hijacked contexts based on the versatile foundation models. We further investigated reforming the hijacked contexts into coherent ones, which promotes future research direction to generate more coherent visual and textual information.

### v2

@@ Abstract
Recently, Large Multi-modal Models (LMMs) have demonstrated their ability to understand the visual contents of images given the instructions regarding the images. Built upon the Large Language Models (LLMs), LMMs also inherit their abilities and characteristics such as in-context learning where a coherent sequence of images and texts are given as the input prompt. However, we identify a new limitation of off-the-shelf LMMs where a small fraction of incoherent images or text descriptions mislead LMMs to only generate biased output about the hijacked context, not the originally intended context. To address this, we investigate whether replacing the hijacked visual and textual contexts with the correlated ones via GPT-4V and text-to-image models can help yield coherent responses.
@@ 1 Introduction
Large Language Models (LLMs) pre-trained with huge amount of text corpus have demonstrated remarkable generalization on various language tasks such as human-like dialogues Google (2023); OpenAI (2023a; b), reasoning-tasks Wei et al. (2021); Lewkowycz et al. (2022); Yao et al. (2022), and few-shot in-context learning Min et al. (2021; 2022) where a few demonstration examples are given as a context for answering the test question. By bootstrapping LLMs with several add-on layers, Large Multi-modal Models (LMMs) OpenAI (2023b); Koh et al. (2023); Liu et al. (2023a); Zhu et al. (2023) were enabled to understand the visual contents and answer the corresponding textual instruction. LMMs also inherited the in-context learning capability of LLM, where a coherent sequence of images and textual information are given as a context for the input prompt Koh et al. (2023); OpenAI (2023b).
In this paper, we identify a new limitation of these off-the-shelf LMMs, hijacking contexts, where a small fraction of incoherent images or text descriptions mislead them to only generate biased output about the incoherent contexts. In Figure 1, we provide the qualitative example of context hijacking on visual story telling (VIST) dataset Ferraro et al. (2016) with FROMAGe Koh et al. (2023) as LMM. Given all the former sequence of images and their corresponding captions, the LMM is asked to provide a caption or description of the final query image. Normally, when all the visual and textual information is coherent and consistently aligned, LMM successfully provides a coherent response under consideration in the given contexts. However, when only a single image-text pair that contains an irrelevant subject is appended to the context, LMM falls into the hijacked context and loses coherency concerning the bunch of formerly given original contexts. In real-world applications and scenarios of LMMs, there is no guarantee for the inexistence of such noises and irrelevant contexts. For reliable usage of LMM, it should be robust to such distribution shift and stick to the majority of context without confusion on the minority of irrelevant contexts.
@@
To address this, we investigate whether modifying the hijacked image and text contexts with the correlated ones can help yield better coherent responses. Firstly, we instructed GPT-4 to identify any irrelevant sentences and replace them with appropriate alternative sentences that convey coherency given in the contexts. Subsequently, we prompted text-to-image diffusion models  [^1] We employed DALLE-3 Betker et al. (2023) for its detailed sensitiveness to the input prompts to generate a coherent image corresponding to the newly replaced sentence under consideration for the original sequence of sentences and images. As a result, a new alternative image and caption that are coherent with the original story plot are interleaved into inputs for LMMs. Notably, we observed that LMMs still exhibit hijacked responses, due to two possible reasons: 1) the generated captions are still confusing which seem to be the combination and mixture between the original and hijacked context. 2) the generated images are not realistic and do not have the consistent texture or styles with the original image sequences.
We showcase our problem statement and aforementioned methods and investigations with qualitative examples one by one [^2] We tested with FROMAGe Koh et al. (2023) as the baseline of LMM.. In summary, our key contributions are outlined as follows:
@@ 2 Related Works
Recently, large language models (LLMs) have gained considerable versatility based on enlarging the scale of pre-training datasets Radford et al. (2019); Brown et al. (2020); OpenAI (2023b). Along with the development of human-aligned feedbacks Christiano et al. (2017), LLMs have further been well-suited for conforming to the instruction rather than predicting simple meaningless next tokens and words. These LLMs have also demonstrated versatility on various language domain tasks including in-context learning Brown et al. (2020); Chan et al. (2022) and long-term dialogues Jozefowicz et al. (2016); Dai et al. (2019); Liu et al. (2023b).
@@
Despite the compelling performance of LLMs on language tasks, LLMs could not understand the visual contexts in an image. To bootstrap LLMs with vision modality, several works Liu et al. (2023a); Zhu et al. (2023); Koh et al. (2023) employed an additional pre-trained visual encoder and leveraged a few linear layers to adapt the visual tokens to be understandable by the frozen LLMs. FROMAGe Koh et al. (2023) further demonstrated the in-context learning capability of LMMs where consecutive sequences of image-text pairs are given as a context to respond to the query image. In this paper, we focus on the vulnerability of LMMs to incoherent and irrelevant context information.
@@
In the language domain, several works empirically observed that LLMs are easily hijacked by irrelevant context information Shi et al. (2023), small perturbations of characters ($e.g.,$ from ”film” to ”fi1m”) Wang et al. (2023), and adversarial addition of suffix words Qiang et al. (2023). In the vision domain, there have been several approaches that investigated the adversarial robustness of LMMs Schlarmann & Hein (2023); Qi et al. (2023) where a small adversarial tweak of pixel values significantly misled LMMs to output violent content. To the best of our knowledge, our work is the first to investigate hijacking contexts in an in-context learning scenario where the consecutive sequence of images and texts along with irrelevant ones are given as the context to provide answers about the query image.
3 Context-Hijacking in Large Multi-modal Models
Preliminary: Large Language Models.
An autoregressive LLM $p_{\theta}$ is pre-trained on a large corpus of text tokens with the next-token prediction objective. Formally, given a text of image caption $y$ and its sequence of tokens ($\bm{s}=s_{1},...,s_{t-1}$) separated by a BPE tokenizer Sennrich et al. (2015), the model is trained to maximize log-likelihood for the next token $s_{t}$:
@@ 3 Context-Hijacking in Large Multi-modal Models
Preliminary: Large Multi-modal Models.
@@
For in-context learning scenarios where multiple images and corresponding text descriptions are given, FROMAGe Koh et al. (2023) is trained with a random concatenation strategy where multiple image-text pairs are piled for attending and responding more explicitly to each image. Hence, FROMAGe exhibited superior capability for answering the query image $x_{n}$ given former visual contexts ($x_{1},\ldots,x_{n-1}$) and textual contexts $(\bm{s}_{1},\ldots,\bm{s}_{n-1})$:
@@ 4 Discussions
We investigate the effect of the sequential location of hijacked context on the coherency of the response from LMMs. In Figure 2, when the hijacking context is interleaved in the former part, the final response still tends to adhere to the original context, ”family”. In contrast, as the hijacking context is interleaved closer to the query image, the response tends to be incoherent to the former contexts and biased towards the hijacked context, ”football”.
@@
Reforming contexts via large foundation models.
Inspired by robustenss of GPT-4V, we investigate whether such large foundation models can identify and reform the irrelevant visual and textual contexts into coherent ones, for better coherent response. As in Figure 4, we first instructed GPT-4 to provide appropriate sentences that convey coherency given in the contexts. Subsequently, we instructed DALLE-3 to generate a synthetic image that corresponds to the newly reformed text, with the consideration of underlying visual and textual context information.
@@
In Figure 3, we observed that our reforming strategy still does not ensure LMMs give coherent responses. We conjecture that the reformed captions still contain the hijacked context (”football”) while also adhering to the original context (”family”). The generated image also correspondingly contains both visual contexts of ”family” and ”football”, confusing LMMs to produce biased responses. We conjecture that the style and texture of the reformed image differ from the original images, which might also induce LMMs to be confused.
@@ 6 Conclusion
In conclusion, we newly identified the context-hijacking problem of existing LMM where the irrelevant visual and textual contexts induce the LMM to generate a biased and incoherent response, disregarding the original context. To mitigate this issue, we introduced a simple method to preemptively remove such hijacked contexts and further reform the hijacked contexts into coherent ones, which promotes future research direction to generate more coherent visual and textual information.
Appendix A Appendix
Figure 4: Reforming hijacked text and images via GPT-4 and DALLE-3. Given multiple image-text pairs of a visual story plot from VIST dataset Ferraro et al. (2016), GPT-4 is instructed to replace any irrelevant images or text descriptions with coherent ones. Subsequently, DALLE-3 is instructed to generate an image corresponding to the newly reformed text, under consideration of the underlying context.